If x : y = 3 : 2 and y : z = 5 : 4, find the ratio of x : y : z
Answer by Boreal(15235)

x/y=3/2
2x=3y, y=(2/3)x
y/z=5/4
4y=5z; y=(5/4)z
Therefore, (2/3)x=(5/4)z
and x=(15/8)z as well as (3/2)y and z=(8/15)x
Therefore, x:(2/3)x:(8/15)x
or using 15 as a common denominator
15:10:8
This would make x:y= 3:2 and y:z=5:4